OpenStudy (anonymous):

Question about finding sum of the arithmetic sequence? So i know the equation is Sn=a(1-r^n)÷(1-r), but usually the common ratio is multiplying by something. Is the formula different when the common ratio is adding?

terenzreignz (terenzreignz):

It's called a common *difference* when arithmetic sequences are involved, and there's a formula for that, too ^_^

OpenStudy (anonymous):

Ohhh okay so i would use a different formula then? thank you!

terenzreignz (terenzreignz):

Yes, you know what it is?

OpenStudy (anonymous):

uhmmm I do not ;o I was going to try googling it but i'm not sure what it would be called? finding the sum of a arithmetic sequence with a common difference? >_<

terenzreignz (terenzreignz):

When an arithmetic sequence is concerned, the sum is given by this formula instead: \[\Large S_n = a_1 + (n-1)d\]where \(a_1\) is the first term \(n\) is the number of terms and \(d\) is the common difference

OpenStudy (anonymous):

okay! i can take it from here, thank you again! ^_^

terenzreignz (terenzreignz):

WAIT NO!

terenzreignz (terenzreignz):

Sorry, my mistake, that's just the nth term. \[\Large a_n = a_1 + (n-1)d\]

OpenStudy (anonymous):

o-o

terenzreignz (terenzreignz):

To get the sum, the formula is

terenzreignz (terenzreignz):

\[\Large S_n = \frac{n(a_1+a_n)}{2}\]

terenzreignz (terenzreignz):

Where n is the number of terms, a1 is the first term, an is the last term,

OpenStudy (anonymous):

hmm well my problem is What is the sum of the arithmetic sequence 3, 9, 15..., if there are 36 terms? so would 15 be the last term? or would i first find the term of 36?

terenzreignz (terenzreignz):

No...the last term is an, find it with this formula\[\Large a_n = a_1 + (n-1)d\]

OpenStudy (anonymous):

okay, thank you again!

terenzreignz (terenzreignz):

Don't thank me until it's over D:

terenzreignz (terenzreignz):

What's the 36'th term?

terenzreignz (terenzreignz):

\[\Large a_n = a_1 + (n-1)d\] in this formula, take n = 36

OpenStudy (anonymous):

uhm i got 213 as the 36th term? ;o

terenzreignz (terenzreignz):

oh okay, good ^_^ \[\large a_{36}=213\] Now find the sum \[\Large S_n = \frac{n(a_1+a_{n})}{2}\] again, n = 36

OpenStudy (anonymous):

I got 3,888 :3

terenzreignz (terenzreignz):

Brilliant. Well done ^_^

OpenStudy (anonymous):

Thank you! ^_^ thank you again for helping me x3

terenzreignz (terenzreignz):

No problem.